What is the ratio of the area of a square to the area of a circle of the same perimeter?

5, A square of length l, a rhombus is out of this triangle making it have an angle of 600 on a single diagonal. What is the area of this rhombus?

6, Solve the inequality tan(x+ ∏/3) ≥ √3

7, Given a triangle with sides 7m, 8m and 3m. What is 50% the area of this triangle?

8, tan (22.5) can be expressed as

A.     (1+20.5)1/2

B.     (-1+20.5)

C.     (2+20.5)1/2

D.     None


1
Expert's answer
2021-09-14T06:09:32-0400

Area of Square "= s \times s = s\\^2"

Area of Circle "= \pi r\\^2"

Perimeter of Square = "4\times s"

Perimeter of Circle = "2\pi r\\"

Perimeter of Square = Perimeter of Circle

"4s = 2 \pi r"

"= s =0.5 \pi r"

"= r = \frac{2s}{\pi}"

Area of Square "= (0.5 \times \pi \times r) \\ ^2 = (0.5 \pi r) \\^2"

Area of circle "= \pi \times (\frac{2s}{\pi})\\^2" "= \frac{4 (s) \\^2}{\pi}"

Substituting "= s =0.5 \pi r" in "= \frac{4 (s) \\^2}{\pi}"

"= \frac{4 (0.5 \pi r) \\^2}{\pi}"

"= \pi r\\^2"

The ratio of the area of a square to the area of a circle is: "= (0.5 \pi r) \\^2 : \pi r\\^2"

"= 0.25 \pi\\^2 r\\^2 : \pi r\\ ^2"

"=0.25 \pi : 1"

5, All squares are rhombuses, but not all rhombuses are squares.

Therefore, the area of the given rhombus is "l \times l" since they have the same area as that of the square.

6,

"\arctan \left(\sqrt{3}\right)+\pi n\le \frac{x+1}{3}<\frac{\pi }{2}+\pi n"

"\arctan \left(\sqrt{3}\right)+\pi n\le \frac{x+1}{3}\quad \mathrm{and}\quad \frac{x+1}{3}<\frac{\pi }{2}+\pi n"

"x\ge \:\pi +3\pi n-1\quad \mathrm{and}\quad \:x<\frac{3\pi -2}{2}+3\pi n"

Answer is: "\pi -1+3\pi n\le \:x<\frac{3\pi -2}{2}+3\pi n"

7) Area of the triangle: "=\sqrt{(s(s-a) (s-b) (s-c)}"

"= s = \frac{p}{2} = \frac{7 + 8 +3}{2} = 9"

"=\sqrt{(9(9-7) (9-8) (9-3)}"

"= \sqrt{108} = 10.39"

50% of the area of the triangle is: "0.5 \times 10.39 = 4.156"


8) The correct answer is: option B :"= (-1 + 2\\ ^\frac{1}{2})"

Explanation:

"=\sqrt{3-2\sqrt{2}}"

"=\sqrt{2}-1"